# Reasoning by Exploration: A Unified Approach to Retrieval and Generation over Graphs

### v1

@@ Abstract.
Reasoning over structured graphs remains a fundamental challenge for Large Language Models (LLMs), particularly when scaling to large graphs. Existing approaches typically follow the retrieval-augmented generation (RAG) paradigm: first retrieving subgraphs relevant to the query and then generating answers conditioned on the retrieved subgraphs. However, such two-phase pipelines often struggle to faithfully incorporate graph structure, since the generation process is ultimately constrained by the quality and completeness of the retrieved subgraph. Although many advanced retrievers have been proposed recently to mitigate this issue, they are usually tailored to the training graphs and generalize poorly to unseen graphs, which limits their practical applicability. In this work, we propose Reasoning by Exploration (RoE), a novel approach that unifies retrieval and generation by framing reasoning over graphs as a process of graph exploration. At each step, the LLM selects candidate nodes and edges to explore, gradually constructing reasoning paths and generating answers along the way. To enable effective exploration, RoE is trained in two stages: supervised fine-tuning (SFT) on gold reasoning paths, followed by reinforcement learning (RL) to enhance exploration effectiveness and generalization. Experiments on benchmark datasets demonstrate that RoE achieves substantial overall improvements over baselines, while also generalizing effectively to unseen graphs.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
@@ 1. Introduction
Large language models (LLMs) (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023) have demonstrated impressive abilities across a wide range of natural language processing tasks. To further enhance their reliability and factuality, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Jiang et al., 2023) has become a widely adopted paradigm. By retrieving external knowledge before generating responses, RAG has achieved strong performance in diverse domains, such as healthcare (Xu et al., 2024), law (Wiratunga et al., 2024), finance (Zhang et al., 2023), and education (Chu et al., 2025). However, most existing RAG systems operate over unstructured text, retrieving passages through lexical or semantic similarity, and therefore do not fully leverage the relational structure inherent in many forms of knowledge.
@@
Graphs provide a natural representation for structured knowledge, capturing rich relational dependencies among entities. They are widely used in domains such as knowledge bases, social networks, and scientific applications including biology and chemistry (Wu et al., 2020; Ma and Tang, 2021; Wu et al., 2023a). Recently, Graph Retrieval-Augmented Generation (GraphRAG) has emerged as a promising way to integrate such graph-structured knowledge into LLMs (Han et al., 2024; Peng et al., 2024). There are mainly two steps in GraphRAG: first, a retriever selects query-relevant content from the entire graph, where the retrieved units may be nodes, paths, or subgraphs; then, a generator, usually an LLM, produces answers conditioned on the retrieved subgraphs. This framework has been successfully applied to tasks such as knowledge graph question answering (KGQA) (Yasunaga et al., 2021, 2022; Zhang et al., 2022a), where answering a query often requires multi-step reasoning over the graphs.
@@
While efficient, this two-phase pipeline may constrain reasoning quality due to the disjoint optimization of retriever and generator. First, the retriever often ignores the specific needs of the generator, which can lead to missing critical evidence or introducing irrelevant noise. For example, consider the multi-hop question “Who are the nephews of Winston Churchill?”. This query may have multiple answers and can be reached through multiple reasoning paths, such as “Churchill $\rightarrow$ sibling $\rightarrow$ child” or “Churchill $\rightarrow$ parent $\rightarrow$ child $\rightarrow$ child”. Heuristic-based retrievers (Yasunaga et al., 2022; Feng et al., 2020) (e.g., retrieving fixed $k$-hop subgraphs) or GNN-based retrievers (Mavromatis and Karypis, 2024; Liu et al., 2024) may include a large amount of irrelevant information about Churchill, overwhelming the generator with noise (Guo et al., 2025a). In contrast, LLM-based retrievers (Luo et al., 2023; Wu et al., 2023b) may generate only a limited set of reasoning paths, leading to missing valid entities and incomplete answers, as shown in Section 3. Second, retrievers or generators trained on specific graphs or datasets usually generalize poorly to unseen graphs as shown in Section 3. For instance, a path-based retriever tuned on a movie knowledge base may learn to generate reasoning paths primarily around relations such as “actor $\rightarrow$ film $\rightarrow$ director”, but it struggles to generalize when applied to graphs with very different relational structures, such as biomedical graphs involving “gene $\rightarrow$ protein $\rightarrow$ disease” paths. This dependence leads to sharp performance drops when deployed on unseen graphs, limiting the robustness and real-world applicability of current GraphRAGs.
@@
To address these limitations, we propose Reasoning by Exploration (RoE), a unified framework that leverages the reasoning ability of LLMs to perform retrieval and generation simultaneously within the same model. Instead of separating the two stages, RoE frames reasoning as a step-by-step graph exploration process, where the model incrementally expands nodes and edges while constructing reasoning paths and generating answers in a unified manner. In this way, RoE has the access to the whole graph. This design directly mitigates the problem of missing or noisy retrieval, since the model actively decides which parts of the whole graph to explore based on its evolving reasoning process, instead of relying on a static retriever. However, it faces two main challenges. First, how can we equip LLMs, which pretrained primarily on unstructured text, with the ability to perform structured graph exploration? Second, how can we ensure that the model learns generalizable exploration strategies rather than memorizing specific graph patterns, so that it can adapt to unseen graphs? To tackle these challenges, we adopt a two-stage training strategy. In the first stage, we construct gold exploration trajectories and apply supervised fine-tuning (SFT) to teach the LLM to expand nodes and edges step by step. In the second stage, we employ reinforcement learning (RL) with different reward signals to further refine the exploration policy, encouraging the model to efficiently discover valid reasoning paths while improving its ability to generalize across diverse and unseen graphs. This combination enables RoE to leverage both explicit supervision and adaptive learning signals, resulting in more faithful reasoning and stronger generalization to unseen graphs. Our key contributions are summarized as follows:
  • We introduce Reasoning by Exploration (RoE), a unified framework that integrates retrieval and generation into a single process by framing reasoning as graph exploration, allowing LLMs to construct reasoning paths and answers simultaneously.
  • We develop a two-stage training strategy that equips LLMs with exploration ability: supervised fine-tuning (SFT) on gold reasoning trajectories to learn step-wise expansion, followed by reinforcement learning (RL) to refine exploration policies and enhance generalization to unseen graphs.
  • We conduct extensive experiments on multi-hop reasoning benchmarks, demonstrating that RoE significantly outperforms state-of-the-art baselines while exhibiting strong generalization across diverse unseen graphs.
@@ 2.1. GraphRAG
GraphRAG (Han et al., 2024; Peng et al., 2024) aims to retrieve relevant information from external graphs to improve the performance and reduce the hallucination of LLMs. Compared to traditional RAG systems (Lewis et al., 2020; Gao et al., 2023) that operate on unstructured text, GraphRAG leverages the relational structure of graphs, where nodes represent entities and edges encode relations between nodes. Similar to RAG, GraphRAG generally consists of two main components: a retriever, which selects query-relevant graph content, and a generator, typically an LLM, which produces answers based on the retrieved subgraphs. To accommodate the graph structure, there are mainly three types of retrievers:
Heuristic-based Retriever. Heuristic-based retrievers typically begin by using similarity-based methods (Wen et al., 2023; Sanmartin, 2024) or entity and relation extraction techniques (Al-Moslmi et al., 2020; Feng et al., 2020) to identify seed nodes and edges that are relevant to the query. These seeds are then expanded by retrieving their $k$-hop neighbors (Yasunaga et al., 2021; Choudhary and Reddy, 2023), forming a candidate subgraph for reasoning. A representative example is G-Retriever (He et al., 2024), which first leverages embedding models to retrieve the most similar nodes and edges to the query, and then applies the Prize-Collecting Steiner Tree (PCST) algorithm to extract a compact subgraph centered around the selected seeds.
GNN-based Retriever. GNN-based retrievers (Liu et al., 2024; Fang et al., 2024) typically train Graph Neural Networks (GNNs) (Wu et al., 2020; Ma and Tang, 2021) to model retrieval as a node classification task, where the correct answer entities are labeled as positive and others as negative. For example, GNN-RAG (Mavromatis and Karypis, 2024) first trains a GNN model and then retrieves the nodes with the highest probability scores as candidate answers, while also identifying the shortest paths connecting the question entities with these candidates to construct supporting evidence. Compared to heuristic-based methods, GNN-based retrievers can capture richer relational dependencies and are more adaptive to complex queries.
@@ 2.2. Reasoning of Large Language Models
Large Language Models (LLMs) have demonstrated strong performance on natural language tasks (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023), but they may struggle with more complex reasoning problems (Zheng et al., 2023; Nezhurina et al., 2024). To further improve their reasoning ability, early works leverage prompting-based methods, such as Chain-of-Thought  (Wei et al., 2022), and Graph-of-Thought  (Besta et al., 2024), which encourage models to generate and organize intermediate reasoning steps during inference. Think-on-graph (ToG) (Sun et al., 2023) also follow this paradigm to prompt LLM reason on the graph. While effective, these methods are often sensitive to prompt design and computationally inefficient. Building on this, more recent approaches (Xu et al., 2025; Sui et al., 2025) focus on supervised fine-tuning (SFT) with curated reasoning traces (Deng et al., 2021; Lobo et al., 2024) and reinforcement learning (RL) with process or outcome rewards (Rafailov et al., 2023; Zhang et al., 2024). These methods allow models to internalize reasoning strategies rather than relying only on inference-time prompting. However, most of these advances focus on text-only tasks, whereas reasoning over graph-structured knowledge introduces additional challenges, requiring models to explore nodes, edges, and multi-hop paths. Although several works (Wang et al., 2023b; Luo et al., 2025; Yu et al., 2025) have attempted to adapt LLMs for graph reasoning tasks, most of them explicitly decompose complex queries into simpler sub-queries and rely on external retrievers. It remains underexplored how to enable LLMs to automatically explore graphs and generate answers in a unified manner.
@@ 3. Preliminary
In this section, we first provide definitions of the key concepts and tasks used throughout the paper. We then analyze the retrievers adopted in current popular GraphRAG methods.
@@ 3.1. Definitions
Knowledge Graphs (KGs). Let $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ denote a knowledge graph, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{R}$ is the set of relations, and $\mathcal{E}$ is the set of edges. Each edge $e\in\mathcal{E}$ represents a directed link between the head and tail entities $h,t\in\mathcal{V}$ under a relation $r\in R$, and can be denoted as a triple $(h,r,t)$. For a given node $v\in\mathcal{V}$, we use $\mathcal{N}_{v}$ to denote its neighboring entities, $\mathcal{R}_{v}$ to denote the associated relations and $\mathcal{E}_{v}$ to denote the connected edges of node $v$.
In this paper, we also augment the knowledge graph with inverse edges. Specifically, for each triple $(h,r,t)\in\mathcal{E}$, we add its inverse edge $(t,r_{\text{inverse}},h)$, following prior works (Kazemi and Poole, 2018; Dutt et al., 2022). This augmentation allows the model to explore the graph in both directions and has been shown to improve reasoning performance.
@@ 3.3. In-Distribution Retriever Performance
In this subsection, we aim to measure the performance of different retrievers when they are both trained and tested on the same dataset (i.e., WebQSP or CWQ). The performance of the three selected methods is shown in Figure 1(a), where we also include the ground-truth Hit and Recall as upper bounds, since the underlying graph does not contain all the answers for every question.
From the results in Figure 1(a), we observe that none of the retrievers achieve perfect performance, even in terms of Hit. For example, the Recall and Precision of G-Retriever on the CWQ dataset are only 38.65 and 5.14, respectively, indicating that many correct answers are missing while substantial noise is included in the retrieved subgraphs. This is because G-Retriever relies on embedding similarity to retrieve nodes, which is often insufficient for handling complex queries. This aligns with the well-known limitation of RAG methods in retrieving evidence for multi-hop queries (Shao et al., 2023; Tang and Yang, 2024), where iterative retrieval strategies are usually adopted by combining intermediate results from the generator. Among the three methods, GNN-RAG performs best, but it still shows around a 10% gap compared to the ground-truth Hit upper bound on the CWQ dataset. This demonstrates that the limitations of current retrievers will substantially constrain downstream generation performance.
@@ 4.1. Exploration Formulation
where each path $p_{i}^{d}$ is defined as
$p_{i}^{d}=[v_{0},r_{1},v_{1},\ldots,r_{k},v_{k}],$
starting from a seed entity $v_{0}\in\mathcal{V}_{0}$ and expanding through a sequence of relations and entities. We define frontier nodes as
@@
The exploration continues until the model predicts that no new paths should be expanded (i.e., $\mathcal{P}_{d+1}=\emptyset$) or a predefined maximum number of steps $D_{\max}$ is reached.
The final outputs include the aggregated predicted answers
$\mathcal{A}=\bigcup_{d}\mathcal{A}_{d},$
which enables the model to find multiple answers through reasoning paths of different lengths.

### v2

@@ Abstract.
Reasoning over structured graphs remains a fundamental challenge for Large Language Models (LLMs), particularly when scaling to large graphs. Existing approaches typically follow the retrieval-augmented generation (RAG) paradigm: first retrieving subgraphs relevant to the query and then generating answers conditioned on the retrieved subgraphs. However, such two-phase pipelines often struggle to faithfully incorporate graph structure, since the generation process is ultimately constrained by the quality and completeness of the retrieved subgraph. Although many advanced retrievers have been proposed recently to mitigate this issue, they are usually tailored to the training graphs and generalize poorly to unseen graphs, which limits their practical applicability. In this work, we propose Reasoning by Exploration (RoE), a novel approach that unifies retrieval and generation by framing reasoning over graphs as a process of graph exploration. At each step, the LLM selects candidate nodes and edges to explore, gradually constructing reasoning paths and generating answers along the way. To enable effective exploration, RoE is trained in two stages: supervised fine-tuning (SFT) on gold reasoning paths, followed by reinforcement learning (RL) to enhance exploration effectiveness and generalization. Experiments on benchmark datasets demonstrate that RoE achieves substantial overall improvements over baselines, while also generalizing effectively to unseen graphs. The code can be found at https://github.com/haoyuhan1/RoE.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792684isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Language modelsccs: Information systems Question answeringccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
Large language models (LLMs) (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023) have demonstrated impressive abilities across a wide range of natural language processing tasks. To further enhance their reliability and factuality, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Jiang et al., 2023; Wang et al., 2024) has become a widely adopted paradigm. By retrieving external knowledge before generating responses, RAG has achieved strong performance in diverse domains, such as healthcare (Xu et al., 2024), law (Wiratunga et al., 2024), finance (Zhang et al., 2023), and education (Chu et al., 2025). However, most existing RAG systems operate over unstructured text, retrieving passages through lexical or semantic similarity, and therefore do not fully leverage the relational structure inherent in many forms of knowledge.
@@
Graphs provide a natural representation for structured knowledge, capturing rich relational dependencies among entities. They are widely used in domains such as knowledge bases, social networks, and scientific applications including biology and chemistry (Wu et al., 2020; Ma and Tang, 2021; Wang et al., 2020). Recently, Graph Retrieval-Augmented Generation (GraphRAG) has emerged as a promising way to integrate such graph-structured knowledge into LLMs (Han et al., 2024; Peng et al., 2024; Lei et al., 2025b). There are mainly two steps in GraphRAG: first, a retriever selects query-relevant content from the entire graph, where the retrieved units may be nodes, paths, or subgraphs; and then, a generator, usually an LLM, produces answers conditioned on the retrieved subgraphs. This framework has been successfully applied to tasks such as knowledge graph question answering (KGQA) (Yasunaga et al., 2021, 2022; Zhang et al., 2022b), where often requires multi-step reasoning over the graphs.
@@
While efficient, this two-phase pipeline may constrain reasoning quality due to the disjoint optimization of retriever and generator. First, the retriever often ignores the specific needs of the generator, which can lead to missing critical evidence or introducing irrelevant noise. For example, consider the multi-hop question “Who are the nephews of Winston Churchill?”. This query may have multiple answers and can be reached through multiple reasoning paths, such as “Churchill $\rightarrow$ sibling $\rightarrow$ child” or “Churchill $\rightarrow$ parent $\rightarrow$ child $\rightarrow$ child”. Heuristic-based retrievers (Yasunaga et al., 2022; Feng et al., 2020) (e.g., retrieving fixed $k$-hop subgraphs) or GNN-based retrievers (Mavromatis and Karypis, 2024; Liu et al., 2024) may include a large amount of irrelevant information about Churchill, overwhelming the generator with noise (Guo et al., 2025b). In contrast, LLM-based retrievers (Luo et al., 2023; Wu et al., 2023) may generate only a limited set of reasoning paths, leading to missing valid entities and incomplete answers, as shown in Section 3. Second, retrievers or generators trained on specific graphs or datasets usually generalize poorly to unseen graphs and question types as shown in Section 3. For instance, a path-based retriever tuned on a movie knowledge base may learn to generate reasoning paths primarily around relations such as “actor $\rightarrow$ film $\rightarrow$ director”, but it may struggle to generalize when applied to graphs in other domains with very different relational structures, such as biomedical graphs involving “gene $\rightarrow$ protein $\rightarrow$ disease” paths. This dependence leads to sharp performance drops when deployed on unseen graphs, limiting the robustness and real-world applicability of current GraphRAGs.
@@
To address these limitations, we propose Reasoning by Exploration (RoE), a unified framework that leverages the reasoning ability of LLMs to perform retrieval and generation simultaneously within the same model. Instead of separating the two stages, RoE frames reasoning as a step-by-step graph exploration process, where the model incrementally expands nodes and edges while constructing reasoning paths and generating answers in a unified manner. In this way, RoE has the access to the whole graph. This design directly mitigates the problem of missing or noisy retrieval, since the model actively decides which parts of the whole graph to explore based on its evolving reasoning process, instead of relying on a static retriever. However, it faces two main challenges. First, how can we equip LLMs, which pretrained primarily on unstructured text, with the ability to perform structured graph exploration? Second, how can we ensure that the model learns generalizable exploration strategies rather than memorizing specific graph patterns, so that it can adapt to unseen graphs? To tackle these challenges, we adopt a two-stage training strategy. In the first stage, we construct gold exploration trajectories and apply supervised fine-tuning (SFT) to teach the LLM to expand nodes and edges step by step. In the second stage, we employ reinforcement learning (RL) to further refine the exploration policy, encouraging the model to efficiently discover valid reasoning paths while improving its ability to generalize across diverse and unseen graphs. Together, this training paradigm enables RoE to combine explicit supervision with adaptive learning signals, leading to more faithful reasoning and stronger generalization. We conduct extensive experiments on multiple multi-hop reasoning benchmarks, showing that RoE consistently outperforms strong state-of-the-art baselines and exhibits robust generalization to unseen graphs and question types across different backbone sizes.
@@ 2.1. GraphRAG
GraphRAG (Han et al., 2024; Peng et al., 2024; Lei et al., 2025a; Han et al., 2025a) aims to retrieve relevant information from external graphs to improve the performance and reduce the hallucination of LLMs. Compared to traditional RAG systems (Lewis et al., 2020; Gao et al., 2023) that operate on unstructured text, GraphRAG leverages the relational structure of graphs, where nodes represent entities and edges encode relations between nodes. Similar to RAG, GraphRAG generally consists of two main components: a retriever, which selects query-relevant graph content, and a generator, typically an LLM, which produces answers based on the retrieved subgraphs. To accommodate the graph structure, there are mainly three types of retrievers:
Heuristic-based Retriever. Heuristic-based retrievers typically begin by using similarity-based methods (Wen et al., 2023; Sanmartin, 2024) or entity and relation extraction techniques (Al-Moslmi et al., 2020; Feng et al., 2020) to identify seed nodes and edges that are relevant to the query. These seeds are then expanded by retrieving their $k$-hop neighbors (Yasunaga et al., 2021; Choudhary and Reddy, 2023), forming a candidate subgraph for reasoning. A representative example is G-Retriever (He et al., 2024), which first leverages embedding models to retrieve the most similar nodes and edges to the query, and then applies the Prize-Collecting Steiner Tree (PCST) algorithm to extract a compact subgraph.
GNN-based Retriever. GNN-based retrievers (Liu et al., 2024; Fang et al., 2024) typically train Graph Neural Networks (GNNs) (Wu et al., 2020; Ma and Tang, 2021) to model retrieval as a node classification task, where the correct answer entities are labeled as positive and others as negative. For example, GNN-RAG (Mavromatis and Karypis, 2024) first trains a GNN model and then retrieves the nodes with the highest probability scores as candidate answers, while also identifying the shortest paths connecting the question entities with these candidates to construct supporting evidence.
@@ 2.2. Reasoning of Large Language Models
Large Language Models (LLMs) have demonstrated strong performance on natural language tasks (Chang et al., 2024; Thirunavukarasu et al., 2023; Kasneci et al., 2023; Han et al., 2025b), but they may struggle with more complex reasoning problems (Zheng et al., 2023; Nezhurina et al., 2024). To further improve their reasoning ability, early works leverage prompting-based methods, such as Chain-of-Thought  (Wei et al., 2022), Tree-of-Thought (Yao et al., 2023) and Graph-of-Thought  (Besta et al., 2024), which encourage models to generate and organize intermediate reasoning steps during inference. Think-on-graph (ToG) (Sun et al., 2023) also follow this paradigm to prompt LLM reason on the graph. While effective, these methods are often sensitive to prompt design and computationally inefficient. Building on this, more recent approaches (Xu et al., 2025; Sui et al., 2025) focus on supervised fine-tuning (SFT) with curated reasoning traces (Deng et al., 2021; Lobo et al., 2024) and reinforcement learning (RL) with process or outcome rewards (Rafailov et al., 2023; Zhang et al., 2024). These methods allow models to internalize reasoning strategies rather than relying only on inference-time prompting. However, most of these advances focus on text-only tasks, whereas reasoning over graph-structured knowledge introduces additional challenges, requiring models to explore nodes, edges, and multi-hop paths. Although several works (Wang et al., 2023a; Luo et al., 2025; Yu et al., 2025) have attempted to adapt LLMs for graph reasoning tasks, most of them explicitly decompose complex queries into simpler sub-queries and rely on external retrievers. It remains underexplored how to enable LLMs to automatically explore graphs and generate answers in a unified manner.
@@ 3. Preliminary
In this section, we first provide definitions, and then analyze the retrievers adopted in current popular GraphRAG methods.
@@ 3.1. Definitions
Knowledge Graphs (KGs). Let $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{E})$ be a knowledge graph, where $\mathcal{V}$ denotes entities, $\mathcal{R}$ relation types, and $\mathcal{E}$ directed edges. Each edge is represented as a triple $(h,r,t)$ with $h,t\in\mathcal{V}$ and $r\in\mathcal{R}$. For any node $v\in\mathcal{V}$, we denote its neighboring entities, incident relations, and connected edges as $\mathcal{N}_{v}$, $\mathcal{R}_{v}$, and $\mathcal{E}_{v}$, respectively. In this paper, we also augment the knowledge graph with inverse edges. Specifically, for each triple $(h,r,t)\in\mathcal{E}$, we add its inverse edge $(t,r_{\text{inverse}},h)$, following prior works (Kazemi and Poole, 2018; Dutt et al., 2022).
@@ 3.3. In-Distribution Retriever Performance
In this subsection, we aim to measure the performance of different retrievers when they are both trained and tested on the same dataset (i.e., WebQSP or CWQ). The performance of the three selected methods is shown in Figure 1(a), where we also include the ground-truth Hit and Recall as upper bounds, since the underlying graph does not contain all the answers for every question. From the results, we observe that none of the retrievers achieve perfect performance. For example, the Recall and Precision of G-Retriever on the CWQ dataset are only 38.65 and 5.14, respectively, indicating that many correct answers are missing while substantial noise is included in the retrieved subgraphs. Among the three methods, GNN-RAG performs best, but it still shows around a 10% gap compared to the ground-truth Hit upper bound on the CWQ dataset. This demonstrates that the limitations of current retrievers will substantially constrain downstream generation performance.
@@ 4.1. Exploration Formulation
where each path $p_{i}^{d}$ is defined as $p_{i}^{d}=[v_{0},r_{1},v_{1},\ldots,r_{k},v_{k}],$ starting from a seed entity $v_{0}\in\mathcal{V}_{0}$ and expanding through a sequence of relations and entities. We define frontier nodes as
@@
The exploration continues until the model predicts that no new paths should be expanded (i.e., $\mathcal{P}_{d+1}=\emptyset$) or a predefined maximum number of steps $D_{\max}$ is reached. The final outputs include all predicted answers $\mathcal{A}=\bigcup_{d}\mathcal{A}_{d}$, which enables the model to find multiple answers through reasoning paths of different lengths.